1. Solve the following: (a) y} – 3y + 34 – y=0. (c) y - y – y + y = 0, y(0) = 2, y(0) = -1, (0) = -2.
Answer #1

(a) yu_34 +34-y=0 91s auxiliary eq. will be m3_ 3m² +3m-1=0 m² 1-3m (m-1) = 0 - (m-1) (m²/1tm) - 3m (m-1) = 0 - (m.1) (m²(c) y_y_yloty=0, Y (o)=2 , Y(o)=1, 4110)=-2. Its auxiliary ean will be m² - m²_m+1=0 or m² (m-1)-1 (m-1)=0 (m-1) (m²-1)=0